OpenStudy (anonymous):

Find the sum of a finite geometric sequence from n = 1 to n = 5, using the expression −3(4)n − 1. -1,023 1,223 1,023 −4,374

OpenStudy (anonymous):

@gswag98

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

hey what class is this

OpenStudy (jdoe0001):

\(\bf -3(4)^{n-1}\) right?

OpenStudy (anonymous):

Yeah c:

OpenStudy (anonymous):

Its Algebra 2 xD

OpenStudy (anonymous):

All i know is, i will never have to use this in life.

OpenStudy (anonymous):

"SNAP" can't help ya then

OpenStudy (jdoe0001):

\(\bf \large { S_{\color{brown}{ n}}=\Sigma_{i=1}^n\ a_1\cdot r^{i-1}=a_1\left(\cfrac{1-r^{\color{brown}{ n}}}{1-r}\right) \\ \quad \\ thus \implies -3(4)^{n-1} \iff a_1\cdot r^{i-1}\to \begin{cases} a_1=-3\\ \bf r=4 \end{cases} \\ \quad \\ S_{\color{brown}{ 5}}=\Sigma_{i=1}^{\color{brown}{ 5}}\ a_1\cdot r^{i-1}=-3\left(\cfrac{1-4^{\color{brown}{ 5}}}{1-4}\right) }\)

OpenStudy (anonymous):

WHAT.

OpenStudy (jdoe0001):

heheh

OpenStudy (jdoe0001):

\(\bf \large S_{\color{brown}{ n}}\) <-- finite sequence sum

OpenStudy (anonymous):

Okay so. 341?

OpenStudy (anonymous):

341*3=1023

OpenStudy (anonymous):

hard 2 understand all that on @jdoe0001 's pic

OpenStudy (anonymous):

OH I GOT IT ;D

OpenStudy (jdoe0001):

well... is a notationi for a geometric sequence =)

OpenStudy (anonymous):

Help with one more? ;3

OpenStudy (jdoe0001):

yeap is 1023